# These $\equalcontThese authors contributed equally to this work.$

### v1

@@ Abstract
Whole slide image (WSI) analysis has emerged as an increasingly essential technique in computational pathology. Recent advances in the pathological foundation models (FMs) have demonstrated significant advantages in deriving meaningful patch-level or slide-level feature representations from WSIs. However, current pathological FMs have exhibited substantial heterogeneity caused by diverse private training datasets and different network architectures. This heterogeneity introduces performance variability when we utilize the extracted features from different FMs in the downstream tasks. To fully explore the advantage of multiple FMs effectively, in this work, we propose a novel framework for the fusion of heterogeneous pathological FMs, called FuseCPath, yielding a model with a superior ensemble performance. The main contributions of our framework can be summarized as follows: (i) To guarantee the representativeness of the training patches, we propose a multi-view clustering-based method to filter out the discriminative patches via multiple FMs’ embeddings. (ii) To effectively fuse the heterogeneous patch-level FMs, we devise a cluster-level re-embedding strategy to online capture patch-level local features. (iii) To effectively fuse the heterogeneous slide-level FMs, we devise a collaborative distillation strategy to explore the connections between slide-level FMs. Extensive experiments conducted on lung cancer, bladder cancer, and colorectal cancer datasets from The Cancer Genome Atlas (TCGA) have demonstrated that the proposed FuseCPath achieves state-of-the-art performance across multiple tasks on these public datasets.
@@
Foundation model, Histopathological image analysis, Multiple instance learning, Multi-model integration, Information fusion
@@ 1 Introduction
Figure 1: (a) Conventional foundation model-based WSI analysis paradigm. To achieve optimal performance on downstream tasks, the most straightforward strategy is to select a patch-level or slide-level foundation model that exhibits the strongest performance on the target task. (b) Proposed heterogeneous foundation model fusion-based WSI analysis paradigm. Based on the concept of ensemble learning, a framework for the simultaneous fusion of heterogeneous patch-level and slide-level pathology FMs will yield a model with superior performance. (c) The key ideas and contributions of this article.
@@
Pathological diagnosis is the gold standard for cancer diagnosis, while whole slide image (WSI) analysis occupies a core position in computational pathology (CPath) and can support key tasks such as cancer subtyping [1, 2], survival prediction [3, 4, 5], and biomarker prediction [6, 7, 8]. In recent years, the rapid development of pathological foundation models (FMs) [9, 10, 11, 12] has brought about a revolutionary transformation in this field.
@@
Current pathology FMs can be categorized into two distinct types, which are patch-level FMs [9, 13, 10] and slide-level FMs [14, 11, 15, 12]. The patch-level FMs are trained with the tiled patches of WSIs. The patch embeddings derived from patch-level FMs will be aggregated with Multiple-Instance Learning (MIL) for the training of downstream tasks in CPath. Most of the patch-level FMs are trained with self-supervised learning methods of different architectures (e.g., Dino-v2 [16] or MAE), using different private datasets. Different from patch-level FMs, the slide-level FMs are capable of constructing slide embeddings with unsupervised learning. Similar to patch-level FMs, the architectures of backbone models and training datasets differ significantly in each slide-level FM. In conclusion, we define these FMs differences as heterogeneity in the pathology FMs.
@@
Because of the heterogeneity, the performance in different downstream tasks and the learned tissue morphologies are diverse across different FMs [17]. To ensure the performance of FMs on downstream tasks, the most trivial strategy is to select a foundation model with the best performance on the corresponding downstream task, as shown in Figure 1 (a). However, this strategy contains obvious shortcomings. Firstly, re-training a foundation model with our own training datasets may not reproduce the optimal performance. Secondly, when we are facing more than one downstream task, re-training many FMs simultaneously is not a flexible solution. Consequently, based on the concept of ensemble learning, it is an effective way to fuse the heterogeneous patch-level and slide-level embeddings from the FMs into a single proxy model for training, as shown in Figure 1 (b). By combining the strengths of each individual foundation model, we will obtain a fused model with improved performance on downstream tasks [17]. However, there still exist two major challenges hindering the fusion of heterogeneous FMs. Firstly, the heterogeneity in the pathology FMs contributes to diverse dimensions and information of the embeddings. It is essential to comprehensively capture the connections between the patch-level or slide-level embeddings derived from heterogeneous FMs. Secondly, the dimensional gaps between patch-level and slide-level embeddings. The representation information captured by patch-level and slide-level embeddings is distributed at different scales. We need to fully leverage the representational information from slide-level embeddings to assist in training models with patch-level embeddings.
@@
To address these challenges, in this work, we propose a novel framework called FuseCPath for the simultaneous fusion of heterogeneous patch-level and slide-level pathology FMs. Figure 1 (c) illustrates the key ideas and contributions of the proposed FuseCPath. The fusion of patch-level FMs aims at fusing the patch-level local features from diverse FMs (Figure 1 (c).1). First, we propose a multi-view clustering strategy to select representative patches utilizing the meaningful features captured from heterogeneous patch-level FMs. Second, we devise a cluster-level feature re-embedding transformer to discover the relations between patch-level FMs in feature space. The fusion of slide-level FMs aims at fusing the slide-level global features from diverse FMs (Figure 1 (c).2). Consequently, we devise a collaborative distillation module to effectively utilize the representative global features residing in slide embeddings as a teacher model. Equipped with the above modules, the FuseCPath framework will be capable of fusing the heterogeneous FMs effectively. The code is publicly available from https://github.com/ZhidongYang/FuseCPath.
@@
  • We propose FuseCPath, a novel framework for fusing the heterogeneous patch-level and slide-level pathology FMs to ensemble a model equipped with better performance. The proposed FuseCPath framework solves the fusion problem by dividing this problem into the following perspectives which are the fusion of patch-level local features from patch-level FMs and the fusion of global features from slide-level FMs.
@@
  • We devise a novel online feature re-embedding transformer that operates on filtered discriminative patch-level features with multi-view clustering. The proposed online re-embedding effectively addresses the issue of fusing the heterogeneous patch FMs by capturing meaningful features locally and connecting the patch embeddings across diverse patch-level FMs.
@@
  • We propose a novel collaborative distillation module for the fusion of slide-level FMs that systematically bridges the gap between the fusion of patch-level and slide-level FMs. The slide-level FMs will serve as teacher models to provide global representations of WSIs.
@@
  • Extensive experiments are done on several public WSI datasets obtained from the Cancer Genome Atlas (TCGA), including lung adenocarcinoma, bladder, and colon adenocarcinoma cancers. The results demonstrate that the FuseCPath will ensemble a new model with superior performance across various downstream tasks, including biomarker prediction, gene expression prediction, and survival analysis.
@@ 2.1 Pathology foundation model
Recent advances in pathology FMs have employed diverse architectural and training paradigms, predominantly utilizing self-supervised learning (SSL) techniques [18, 19] to extract meaningful representations from unannotated patches in WSIs. SSL-based approaches can be summarized as follows: (1) contrastive learning frameworks such as REMEDIS [20], which adapt SimCLR frameworks [19] by maximizing feature similarity between comparable regions within individual WSIs while minimizing the similarity across disparate slide regions; (2) masked image modeling adopted by BEPH [21], Prov-gigapath [14] and CONCH [9], where random patch occlusion forces models to reconstruct masked tissue patterns, thereby capturing robust contextual relationships; and (3) knowledge distillation implementations exemplified by Virchow2 [10], UNI [13], and Hibou [22], which employ DINO-based frameworks [16] to transfer knowledge from teacher to student models, yielding compact yet generalizable representations without extensive annotations.
@@
2.2 Multiple instance learning
@@
Multi-instance learning is a widely adopted weakly supervised learning strategy in the applications of downstream tasks for WSI analysis [28, 29, 30] due to the lack of annotations. The attention-based deep multi-instance learning (AB-MIL) proposed by [31] first adopts convolutional neural networks (CNNs) to multi-instance learning. This technique is widely extended to the application for WSI image analysis. [28] introduced a recalibrated multi-instance learning framework (RMDL) for the classification of whole slide images (WSIs) of gastric tissues. The RMDL employs a convolutional neural network (CNN) to identify discriminative instances (the image patches) within each WSI and subsequently trains the model exclusively on these selected instances. RMDL captures dependencies among instances and dynamically recalibrates their features based on the coefficients derived from fused feature representations. [30] developed a dual-stream multiple instance learning network (DSMIL) comprising two synergistic streams: one learns an instance-level classifier using max-pooling to identify the highest-scoring (critical) instance, while the other computes attention scores for instances based on their proximity to the critical instance. [32] proposed DeepAttnMISL, a survival prediction model that integrates attention mechanisms with multi-instance learning. This approach clusters the patches extracted from WSIs into phenotypically distinct groups, selects representative patches from each cluster, and processes them through a Siamese multi-instance fully convolutional network. The model subsequently aggregates features via attention-based multiple instance learning (AB-MIL) pooling to predict patient survival risk. Similarly, [29] proposed CLAM, which also operates in two stages: first, patches are encoded into feature vectors using a pre-trained CNN, and then these features are processed by a clustering-constrained attention mechanism within a multiple instance learning framework to produce final predictions. [33] developed DeepSMILE, a two-stage framework wherein the first stage employs the contrastive learning method SimCLR for patch-level feature extraction, generating representative feature embeddings. The second stage incorporates these features into the proposed VarMIL, which is an extension of AB-MIL that introduces a feature variability module to explicitly model tumor heterogeneity. Yan et al. [6] proposed a hierarchical deep multi-instance learning-based framework called HD-MIL to accurately predict gene mutations in bladder cancer by leveraging a contrastive learning framework called Bootstrap Your Own Latent (BYOL) to derive high-quality feature representations. Yang et al. [34] proposed to incorporate the Selective Scan Space State Sequential Model (Mamba) in Multiple Instance Learning (MIL) for long sequence modeling with linear complexity to adjust the high-resolution of WSIs. Similarly, Tang et al. [35] proposed a re-embedding strategy called R2T to online captures foundation model-level local features and establishes connections across different regions. Additionally, the proposed R2T can be integrated into MIL models (named as R2T-MIL) to improve the performance of several downstream tasks.
@@ 2.3 Patch selection in WSI analysis
The clustering strategy is also an effective way to provide prior knowledge for patch selection. Based on the result of the feature clustering, several clustering-based methods ([42, 43, 6]) are proposed to guarantee the representativeness of the selected patches. The survival prediction method (WSISA) developed by [42] can make effective use of all distinguishing patch features in WSIs, thereby significantly enhancing survival prediction performance compared to existing methods. WSISA first selects hundreds of patches from each WSI and then further clusters these selected patches. Then, it selects clusters based on the patch-level prediction performance using CNN, and combines the chosen clusters to make the final prediction. Based on data from 253 bladder cancer patients in the TCGA dataset, the method proposed by [43] firstly combines the advantages of selecting patches in Regions-of-Interest (RoI) from detected cancer areas and clusters. [6] proposed to select representative patches from clustered detected cancer areas using high-quality embeddings derived from a BYOL-based pre-trained model.
@@ 3 Method
Figure 2: The overall architecture and main components of the proposed FuseCPath framework. (a) The overall architecture of the FuseCPath framework. The FuseCPath can be divided into two essential branches, which are patch-level features re-embedding and slide-level features collaborative distillation. (b) The demonstration of patch-level features re-embedding. Representative features can be summarized with clustering and sparse re-embedding. (c) (c) The re-embedded features aggregation module is implemented by AB-MIL.
FuseCPath is a framework for the fusion of heterogeneous pathological FMs, which contributes to a significant performance improvement on the WSI image analysis. We provide a brief overview of FuseCPath below, and the Figure 2 presents more details. Given a set of WSIs $\{X_{i}|X_{i}\in\mathbb{R}^{d_{x}\times d_{y}\times 3}\}$, FuseCPath will simultaneously process the patch embeddings and slide embeddings of $X_{i}$. FuseCPath will first extract the embeddings with multiple patch FMs and slide FMs. For the fusion of patch-level FMs, the FuseCPath will first cluster the patch embeddings with multiview spectral clustering to find representative patches. Then, a Cluster-level Re-embedding Transformer (CR2T) is used to online fuse the patch embeddings, and the Attention-Based Multiple instance learning (AB-MIL) to aggregate the re-embedded features. For the fusion of slide-level FMs, FuseCPath regards the slide embeddings as the teacher models’ information, with a collaborative distillation loss for the model training.
@@ 3.1 Multi-view patch features clustering
Due to the extremely high resolution of the WSIs, it will be a challenging operation to input all the patches extracted from the WSIs for training. Consequently, a typical solution is to select a subset of the patches randomly with a fixed amount of patches. However, the random selection can not guarantee the representativeness of the patches for training. In this work, we devise a cluster-based strategy to select representative patch embeddings for training.
@@ 3.2 Patch-level features re-embedding
where $z$ denotes the aggregated slide-level features using patch-level features from multiple sources of FMs $f^{i}_{pe}$. $\mathcal{A}(\cdot)$ denotes the mapping function of feature aggregation. The performance of this strategy is limited by the difference between our own fine-tuning datasets and the original training datasets for the FMs. A more effective solution is based on the online simultaneous training using a consistent training paradigm. Thus, we devise an online patch-level features re-embedding strategy to fuse the embeddings from heterogeneous FMs into a single model. This strategy can be formulated as follows.
@@
where $\mathcal{R}(\cdot)$ denotes the mapping function of online features re-embedding. Inspired by R2Transformer in [35], we opt for the Regional Multi-head Self-attention (R-MSA) and Cross Regional Multi-head Self-Attention (CR-MSA) as the base model for our strategy. Figure 1 (b) demonstrates the procedure of online re-embedding in our FuseCPath.
@@
$\hat{\mathbf{Z}}^{pe}_{m}=\text{MSA}(\text{LN}([\mathbf{H}^{f^{1}_{pe}}_{s},.,\mathbf{H}^{f^{N}_{pe}}_{s}]))+$
$[\mathbf{H}^{f^{1}_{pe}}_{s},.,\mathbf{H}^{f^{N}_{pe}}_{s}],\hat{\mathbf{Z}}^{pe}\in\mathbb{R}^{(N_{K}\times K)\times D}$

### v2

@@ Abstract
Whole slide image (WSI) analysis has emerged as an increasingly essential technique in computational pathology. Recent advances in the pathology foundation models (FMs) have demonstrated significant advantages in deriving meaningful patch-level or slide-level multi-scale features from WSIs. However, current pathology FMs have exhibited substantial heterogeneity caused by diverse private training datasets and different network architectures. This heterogeneity introduces performance variability when we utilize the features from different FMs in the downstream tasks. To fully explore the advantages of multiple FMs effectively, in this work, we propose a novel framework for the fusion of multi-scale heterogeneous pathology FMs, called FuseCPath, yielding a model with a superior ensemble performance. The main contributions of our framework can be summarized as follows: (i) To guarantee the representativeness of the training patches, we propose a multi-view clustering-based method to filter out the discriminative patches via multiple FMs’ embeddings. (ii) To effectively fuse the patch-level FMs, we devise a cluster-level re-embedding strategy to online capture patch-level local features. (iii) To effectively fuse the slide-level FMs, we devise a collaborative distillation strategy to explore the connections between slide-level FMs. Extensive experiments demonstrate that the proposed FuseCPath achieves state-of-the-art performance across multiple tasks on diverse datasets.
@@
Foundation model, Histopathological image analysis, Multi-model integration, Information fusion
@@ 1 Introduction
Figure 1: (a) Conventional foundation model-based WSI analysis paradigm. To achieve optimal performance on downstream tasks, the most straightforward strategy is to select a patch-level or slide-level foundation model that exhibits the strongest performance on the target task. (b) Proposed heterogeneous foundation model fusion-based WSI analysis paradigm. Based on the concept of ensemble learning, a framework for the fusion of multi-scale (patch-level and slide-level) heterogeneous pathology FMs will yield a model with superior performance. (c) The key ideas and contributions of this article.
@@
Pathological diagnosis is the gold standard for cancer diagnosis, while whole slide image (WSI) analysis occupies a core position in computational pathology (CPath) and can support key tasks such as cancer subtyping [1, 2], survival prediction [3, 4, 5], and biomarker prediction [6, 7, 8]. In recent years, the rapid development of pathology foundation models (FMs) [9, 10, 11, 12] has brought about a revolutionary transformation in this field.
@@
Current pathology FMs can be categorized into two distinct types at different scales, which are patch-level FMs [9, 13, 10] and slide-level FMs [14, 11, 15, 12]. The patch-level FMs are trained with the tiled patches of WSIs. The patch embeddings derived from patch-level FMs will be aggregated with Multiple-Instance Learning (MIL) for the training of downstream tasks in CPath. Most of the patch-level FMs are trained with self-supervised learning methods of different architectures (e.g., Dino-v2 [16] or MAE [17]), using different private datasets. Different from patch-level FMs, the slide-level FMs are capable of constructing slide embeddings with unsupervised learning. Similar to patch-level FMs, the architectures of backbone models and training datasets differ significantly in each slide-level FM. In conclusion, we define these FMs differences as heterogeneity in the pathology FMs.
@@
Because of the heterogeneity, the performance in different downstream tasks and the learned tissue morphologies are diverse across different FMs [18]. To ensure the performance of FMs on downstream tasks, the most trivial strategy is to select a foundation model with the best performance on the corresponding downstream task, as shown in Figure 1(a). However, this strategy contains obvious shortcomings. Firstly, re-training a foundation model with our own training datasets may not reproduce the optimal performance. Secondly, when we are facing more than one downstream task, re-training many FMs simultaneously is not a flexible solution. Consequently, based on the concept of ensemble learning, it is an effective way to fuse the patch-level and slide-level embeddings from the heterogeneous FMs into a single proxy model for training, as shown in Figure 1(b). By combining the strengths of each individual foundation model, we will obtain a fused model with improved performance on downstream tasks [18]. However, there still exist two major challenges hindering the fusion of heterogeneous FMs. Firstly, the heterogeneity in the pathology FMs contributes to diverse dimensions and information of the embeddings. It is essential to comprehensively capture the connections between the patch-level or slide-level embeddings derived from heterogeneous FMs. Secondly, the scale gaps between patch-level and slide-level embeddings. The representation information captured by patch-level and slide-level embeddings is distributed at different scales. We need to fully leverage the representational information from slide-level embeddings to assist in training models with patch-level embeddings.
@@
To address these challenges, in this work, we propose a novel framework called FuseCPath for the fusion of multi-scale (patch-level and slide-level) heterogeneous pathology FMs. Figure 1(c) illustrates the key ideas and contributions of the proposed FuseCPath. The fusion of patch-level FMs aims at fusing the patch-level local features from diverse FMs (Figure 1(c)). First, we propose a multi-view clustering strategy to select representative patches utilizing the meaningful features captured from heterogeneous patch-level FMs. Second, we devise a cluster-level feature re-embedding transformer to discover the relations between patch-level FMs in feature space. The fusion of slide-level FMs aims at fusing the slide-level global features from diverse FMs (Figure 1(c)). Consequently, we devise a collaborative distillation module to effectively utilize the representative global features residing in slide embeddings as a teacher model. Equipped with the above modules, the FuseCPath framework will be capable of fusing the heterogeneous FMs effectively. The code is publicly available from https://github.com/ZhidongYang/FuseCPath.
@@
  • We propose FuseCPath, a novel framework for fusing the heterogeneous multi-scale (patch-level and slide-level) pathology FMs to integrate a model with better performance.
@@
  • We devise a novel online feature re-embedding transformer that operates on filtered discriminative patch-level features with multi-view clustering. The proposed online re-embedding effectively addresses the issue of fusing the heterogeneous patch-level FMs by capturing meaningful features locally and connecting the patch embeddings across diverse patch-level FMs.
@@
  • We propose a novel collaborative distillation module for the fusion of slide-level FMs that systematically bridges the scale gap between the fusion of patch-level and slide-level FMs. The slide-level FMs will serve as teacher models to provide global representations of WSIs.
@@
  • Extensive experiments demonstrate that the FuseCPath will ensemble a new model with superior performance across various downstream tasks in WSI analysis.
@@ 2.1 Pathology foundation model
Recent advances in pathology FMs have employed diverse architectural and training paradigms, predominantly utilizing self-supervised learning (SSL) techniques [19, 20] to extract meaningful representations from unannotated patches in WSIs. SSL-based approaches can be summarized as follows: (1) contrastive learning frameworks such as REMEDIS [21], which adapt SimCLR frameworks [20] by maximizing feature similarity between comparable regions within individual WSIs while minimizing the similarity across disparate slide regions; (2) masked image modeling adopted by Prov-gigapath [14], CONCH [9], and BEPH [22], where random patch occlusion forces models to reconstruct masked tissue patterns, thereby capturing robust contextual relationships; and (3) knowledge distillation implementations exemplified by Virchow2 [10], UNI [13], and Hibou [23], which employ DINO-based frameworks [16] to distill the knowledge from teacher to student models, yielding compact yet generalizable representations without extensive annotations.
@@
2.2 Multiple instance learning in WSI analysis
@@
Multi-instance learning (MIL) is a predominant weakly supervised learning strategy widely adopted in the applications of downstream tasks for WSI analysis [29, 30, 31], which solves the problem of lacking high-quality annotations. The attention-based deep multi-instance learning (AB-MIL) proposed by [32] first adopts convolutional neural networks (CNNs) to multi-instance learning. This technique is widely extended to the application for WSI image analysis. [29] introduced a recalibrated multi-instance learning framework (RMDL) for the classification of whole slide images (WSIs) of gastric tissues. The RMDL employs a convolutional neural network (CNN) to identify discriminative instances (the image patches) within each WSI and subsequently trains the model exclusively on these selected instances. RMDL captures dependencies among instances and dynamically recalibrates their features based on the coefficients derived from fused feature representations. [31] developed a dual-stream multiple instance learning network (DSMIL) comprising two synergistic streams: one learns an instance-level classifier using max-pooling to identify the highest-scoring (critical) instance, while the other computes attention scores for instances based on their proximity to the critical instance. [33] proposed DeepAttnMISL, a survival prediction model that integrates attention mechanisms with multi-instance learning. This approach clusters the patches extracted from WSIs into phenotypically distinct groups. Then it selects representative patches from each cluster and processes them through a Siamese multi-instance fully convolutional network. The model subsequently aggregates features via attention-based multiple instance learning (AB-MIL) pooling to predict patient survival risk. Similarly, [30] proposed CLAM, which also operates in two stages: first, patches are encoded into feature vectors using a pre-trained CNN, and then these features are processed by a clustering-constrained attention mechanism within a multiple instance learning framework to produce final predictions. [34] developed DeepSMILE, a two-stage framework wherein the first stage employs the contrastive learning method SimCLR for patch-level feature extraction, generating representative feature embeddings. The second stage incorporates these features into the proposed VarMIL, which is an extension of AB-MIL that introduces a feature variability module to explicitly model tumor heterogeneity. Yan et al. [6] proposed a hierarchical deep multi-instance learning-based framework called HD-MIL to accurately predict gene mutations in bladder cancer by leveraging a contrastive learning framework called Bootstrap Your Own Latent (BYOL) to derive high-quality feature representations. Yang et al. [35] proposed to incorporate the Selective Scan Space State Sequential Model (Mamba) in Multiple Instance Learning (MIL) for long sequence modeling with linear complexity to adjust the high-resolution of WSIs. Similarly, Tang et al. [36] proposed a re-embedding strategy called R2T to online captures foundation model-level local features and establishes connections across different regions. Additionally, the proposed R2T can be integrated into MIL models (R2T-MIL) to improve the performance of several downstream tasks.
@@ 2.3 Patch selection in WSI analysis
The clustering strategy is also an effective way to provide prior knowledge for patch selection. Based on the result of the feature clustering, several clustering-based methods ([43, 44, 6]) are proposed to guarantee the representativeness of the selected patches. The survival prediction method (WSISA) developed by [43] can make effective use of all distinguishing patch features in WSIs, thereby significantly enhancing survival prediction performance compared to existing methods. WSISA first selects hundreds of patches from each WSI and then further clusters these selected patches. Then, it selects clusters based on the patch-level prediction performance using CNN, and fuses them to make the final prediction. Inspired by this method, [44] firstly combines the advantages of selecting patches in Regions-of-Interest (RoI) from detected cancer areas and clusters, using the data from 253 bladder cancer patients in the TCGA dataset. The method proposed by [6] proposed to select representative patches from clustered detected cancer areas using high-quality embeddings derived from a BYOL-based pre-trained model.
@@ 3 Method
Figure 2: The overall architecture and main components of the proposed FuseCPath framework. (a) The overall architecture of the FuseCPath framework. The FuseCPath can be divided into two essential branches, which are patch-level features re-embedding and slide-level features collaborative distillation. (b) The demonstration of patch-level features re-embedding. Representative features can be summarized with multi-view clustering and online re-embedding. (c) The re-embedded features aggregation module is implemented by AB-MIL.
FuseCPath is a framework for the fusion of multi-scale (patch-level and slide-level) heterogeneous pathology FMs, which contributes to a significant performance improvement on the WSI image analysis across multiple tasks. We provide a brief overview of FuseCPath below, and Figure 2(a) presents more details. Given a set of WSIs $\{X_{i}|X_{i}\in\mathbb{R}^{d_{x}\times d_{y}\times 3}\}$, FuseCPath will simultaneously fuse the patch embeddings and slide embeddings of $X_{i}$. For the fusion of patch-level FMs, the FuseCPath will first cluster the patch embeddings with multiview spectral clustering to find representative patches. Then, a Cluster-level Re-embedding Transformer (CR2T) is used to online fuse the patch embeddings, and the Attention-Based Multiple instance learning (AB-MIL) to aggregate the re-embedded features. For the fusion of slide-level FMs, FuseCPath regards the slide-level embeddings as the teacher models’ information, which is implemented with a collaborative distillation during training.
@@ 3.1 Multi-view patch features clustering
Due to the extremely high resolution of the WSIs, it will be a challenging operation to input all the patches extracted from the WSIs for training. Consequently, a typical solution is to select a subset of the patches randomly with a fixed amount of patches. However, the random selection can not guarantee the representativeness of the patches for training. In this work, we devise a cluster-based strategy to select representative patch embeddings from multiple FMs for training.
@@ 3.2 Patch-level features re-embedding
where $z$ denotes the aggregated slide-level features using patch-level features from multiple sources of FMs $f^{i}_{pe}$. $\mathcal{A}(\cdot)$ denotes the mapping function of feature aggregation. The performance of this strategy is limited by the difference between our own fine-tuning datasets and the original training datasets for the FMs. A more effective solution is based on online simultaneous training using a consistent training paradigm. Thus, we devise an online cluster-level patch features re-embedding strategy to fuse the embeddings from heterogeneous FMs into a single model. This strategy can be formulated as follows.
@@
where $\mathcal{R}(\cdot)$ denotes the mapping function of online features re-embedding. Inspired by R2Transformer in [36], we opt for the Regional Multi-head Self-attention (R-MSA) and Cross Regional Multi-head Self-Attention (CR-MSA) as the base model for our strategy. Figure 2(b) demonstrates the procedure of online re-embedding for patch-level features in our FuseCPath.
@@
$\hat{\mathbf{Z}}^{pe}_{m}=\text{MSA}(\text{LN}([\mathbf{H}^{f^{1}_{pe}}_{s},.,\mathbf{H}^{f^{N}_{pe}}_{s}]))+[\mathbf{H}^{f^{1}_{pe}}_{s},.,\mathbf{H}^{f^{N}_{pe}}_{s}],$
$\hat{\mathbf{Z}}_{m}^{pe}\in\mathbb{R}^{(N_{K}\times K)\times D},$